OpenStudy (anonymous):

multiply 4xy3 over x2y * y over 8x then simplify result

OpenStudy (anonymous):

\[\frac{4xy^3}{x^2y}\times \frac{y}{8x}\]?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[\frac{4xy^3}{8x^3y}\]then cancel

OpenStudy (anonymous):

\[=\frac{y^2}{2x^2}\]